# A Low Rank Neural Representation of Entropy Solutions

### v1

@@ 1 Introduction
The limitations of classical reduced models is readily seen when one attempts to produce one for the simplest of hyperbolic problems, even the advection equation [73, 15]. The cause is the slow decay rate of the Kolmogorov width [63] of the solution manifold; its lower bounds imply that there cannot be a small set of reduced basis functions whose superpositions can approximate the solution uniformly well [82, 59, 30]. More generally, non-classical reduced models are also not completely immune from such lower bounds; slow decay can be established even for nonlinear benchmarks such as the stable width and entropy for high-dimensional parametric transport equations [14, 72]. Also relevant in this context, there are studies of stability of classical reduced models [31, 57, 20, 1].
Various attempts have been made to obtain the reduced complexity for convection dominated problems by incorporating the structure of the transport. To provide a possibly incomplete list of proposed approaches that are the most relevant to this work: template-fitting [73], method of freezing [4, 58], shock reconstruction [15], approximated Lax-Pairs [27], advection modes [39], transported snapshot interpolation (TSI) [82], shifted proper orthogonal decomposition (sPOD) [66], calibrated manifolds [9], Lagrangian basis method [53], transport reversal [69], registration methods [77], Wasserstein barycenters [23] or quadratic manifolds [25]. Some of these proposed methods can be viewed as dynamical low-rank approximations [42, 74].
@@
The trained LRNRs can serve as a reduced model approximation of the family of solutions to the parametrized PDEs: A rapid approximation is achieved due to the small degrees of freedom, much like the rapid computations performed during the online phase for classical reduced models. Moreover, it is not difficult to see that a reduction in online complexity can be achieved by utilizing the low rank structure in the weights and biases, much like the classical reduced models [37]. This is a property LRNR inherits by being a direct generalization of the compositional reduced model utilizing transported subspaces [70].
@@ 2.1 Entropy solutions
Characteristic curves of the initial value problem Eq. 2.1 are given by the solutions $X:\Omega_{x}\times\Omega_{t}\to\mathbb{R}$ to the family of ordinary differential equations
@@
This formulation for the characteristic curves is also called generalized characteristics [18]. It is straightforward to see that the problem Eq. 2.5 has a unique solution. This is due to the following: The RHS $G$ is locally continuous, so the unique solution exists locally where it is continuous [32]; the solution is uniformly stable in neighborhoods of the points of discontinuity due to the entropy conditions, and at points of discontinuity the shock curve is also uniquely determined by the Rankine-Hugoniot jump conditions [75]. One also observes that the unique solution $X$ is a continuous function of the initial value $x$, as $G$ is of bounded variation.
@@
We refer to both functions Eq. 2.5 and Eq. 2.6 as $G$ and distinguish them by the number of the independent variables (i.e. number of input arguments).
Since our problem has no source terms Eq. 2.1, it is constant along the characteristic curves. So we may rewrite Eq. 2.5 as follows,
@@ 2.2 Rarefied characteristics
In this section, we modify the differential equation describing the characteristic curves Eq. 2.7, so that the form of the entropy solution Eq. 2.11 can be simplified. The modified form of the solution combines the two cases into just one. We shall achieve this by extending the initial condition $u_{0}$ and the characteristic curves $X$ to an extended spatial domain that is larger than the original spatial domain $\Omega_{x}$.
@@
Figure 2: An example of an extended initial condition $\hat{u}_{0}$ of $u$.
@@
and $\widehat{G}:\Omega_{x}\times\widehat{\Omega}_{x}\times\Omega_{t}\to\mathbb{R}$ is an extension of $G$ Eq. 2.6,
@@
The nonlinear RHS term Eq. 2.24 of the ODE Eq. 2.22 can be written as a function of bounded variation that depends only on $\widehat{X}$ and $t$. This implies that $\widehat{X}(\cdot,t)$ is non-decreasing for all $t\in\Omega_{t}$, regardless of whether $u_{0}$ has jumps (see, e.g. Ch. 5 of [32]); positive jumps of $u_{0}$ are replaced by continuous pieces in $\widehat{\Omega}_{x}$, negative jumps are replaced by piecewise constants.
See Figure 3 for an illustration showing $\hat{I}$ and its role in relating $u_{0}$ with $\hat{u}_{0}$ Equation 2.22. A diagram depicting the evolution of $\widehat{X}$ is shown in Figure 4.
Figure 3: An example of the extension $u_{0}$ to $\hat{u}_{0}$ as in Eq. 2.20 in the case of the Burgers flux $F(u)=u^{2}/2$.
@@
The entropy solution $u$ to Eq. 2.1 is given by
@@
Recall that $(\cdot)^{+}$ denotes the left-inverse Eq. 2.10.
@@
(2.26) $\widehat{X}^{+}(x,t)=\iota\circ X^{+}(x,t)\quad\text{ for }(x,t)\in\Omega\smallsetminus\mathcal{R}.$
Otherwise if $(x,t)\in\mathcal{R}$, the expression yields the similarity solution: suppose $y\in\widehat{\Omega}_{x}$ satisfies $x=\hat{I}(y)+tF^{\prime}\circ\hat{u}_{0}(y)$ then one can show that $(x-\bar{\xi}(x,t))/t=F^{\prime}\circ\hat{u}_{0}(y)$, implying $\hat{u}_{0}(y)=(F^{\prime})^{-1}[(x-\bar{\xi}(x,t))/t]$. This agrees with the entropy solution Eq. 2.11.
We make a couple of remarks here. Note that new shocks cannot form at $\widehat{X}(x_{0},t)$ for $x_{0}\in\Gamma_{x}$ since the solution to Eq. 2.22 is strictly increasing with respect to $x$ at $x_{0}$. So the solution given by Eq. 2.25 at the points $(\widehat{X}(x_{0},t),t)$, for times before $\widehat{X}(x_{0},\cdot)$ enters a shock, agrees with the expression for the rarefaction fan. These observations imply that $\widehat{X}$ is continuous.
@@ 2.4 Compositional form
In the previous section, we have simplified the expression of the entropy solution from Eq. 2.11 to Eq. 2.26, by incorporating the evolution of the rarefaction waves into the characteristic curves in Eq. 2.22. Next, we simplify how the shock is represented in $\widehat{X}$.
In the classical regime, the characteristic curves Eq. 2.5 are represented by a linear superposition,
@@
Figure 5: A stationary shock example. Upon shock formation, a kink appears in $\widehat{X}$ and travels to the left and to the right.
@@
In this representation, the time-dependence is in the coefficients of linear combinations within each composition, and the dependence of the coefficients with respect to $t$ is linear, in a similar manner as the classical case.
@@
Figure 6: Relief characteristic curves after shock formation time $t_{S}$.
@@ 3 Low Rank Neural Representation
We introduce a family of feedforward neural networks that satisfy a low rank condition. The condition is inspired by the fact that the entropy solution, represented using the relief characteristic curves Eq. 2.35, is made up of compositions of low rank representations. To make this point clear, we draw up a diagram of each step of the evaluation of Eq. 2.35 along with the associated linear spaces, as follows.
@@
For example, the operations in the second step can be written in the low rank form
@@
where $z=(x,t\varphi_{1}(x)-\varphi_{2}(x))^{T}$ is the output from the previous step. This expression has the appearance of a vector with a low rank representation, if one views the function values $(\phi_{i}(z))$ as column vectors of infinite length.
This structure closely resembles that of a neural network. The only component that requires a close scrutiny is the left-inverse $(\cdot)^{+}$ and its treatment will be a key discussion in the coming sections (Section 4.1 and Section 5.1). For the moment, let us focus on establishing an analogy with feedforward neural networks. We begin by introducing notations and definitions of a neural network. Our notations are similar to those in common use (see, e.g. [21, 3], for similar definitions). Throughout, the superscript $\ell$ will index the layer number.
@@
in which $\odot$ is the entrywise composition, and the affine maps $A_{\ell}$ are given by
@@
Observe how one can view the superpositions in between the compositions in Eq. 3.1 as being analogous to the affine mappings that appear in the neural network Eq. 3.6. With this viewpoint, it is natural to consider a neural network with the affine mappings with low rank weights and biases, in correspondence to the low rank representation in Eq. 3.1. That is, in studying entropy solutions, it is natural to turn our attention to neural networks whose weights $W$ and biases in $B$ are represented as linear combinations of a few rank-1 linear transformations. For example, we can require that
(3.7) $W^{\ell}=\sum_{i=1}^{r}\gamma_{i}^{\ell}U_{i}^{\ell},\quad B^{\ell}=\sum_{i=1}^{r}\theta_{i}^{\ell}V_{i}^{\ell},$
@@
The coefficients $\{\gamma^{\ell}_{i},\theta^{\ell}_{i}\}$ correspond to the coefficients $\{1,t\}$ used in the linear combinations at each compositional step in Eq. 3.1. One notes that the latter coefficients depend on the parameter $t$ (the time variable), and the role of this parameter is different from either the input $x$ or the fixed weights and biases $\{U_{i}^{\ell},V_{i}^{\ell}\}$ prescribing the linear spaces the weights and biases belong to. It is therefore natural to view the coefficients $\{\gamma_{i}^{\ell},\theta_{i}^{\ell}\}$ as distinct external parameters; a specific set of coefficients $\{\gamma_{i}^{\ell},\theta_{i}^{\ell}\}$ correspond to a specific neural network belonging to the family of neural networks whose weights and biases satisfy Eq. 3.7. The dependence on $t$ in Eq. 3.1 would correspond to the dependence of the coefficients $\{\gamma_{i}^{\ell},\theta_{i}^{\ell}\}$ on $t$. On the other hand, the fixed weights and biases correspond to the functions in the low rank representation that do not depend on $t$, e.g. the functions $\{\varphi_{i},\phi_{j},\psi_{k}\}$ in Eq. 3.2.
@@
In summary, a LRNR is a family of feedforward networks that have weights and bias at each individual layer that belong to a fixed linear subspace of dimension at most $r$. The aim of the following sections is to establish the connection between LRNR (Definition 3.3) and the entropy solution Eq. 3.1. Throughout this work, our focus is on scalar problems in a single spatial dimension, so the input and output dimensions are one, that is, $M_{0}=M_{L}=1$.

### v2

# explExample

@@ 1 Introduction
The limitations of classical reduced models are readily seen when one attempts to produce one for the simplest of hyperbolic problems, even the advection equation [74, 16]. The cause is the slow decay rate of the Kolmogorov width [64] of the solution manifold; its lower bounds imply that there cannot be a small set of reduced basis functions whose superpositions can approximate the solution uniformly well [83, 60, 31]. More generally, non-classical reduced models are also not completely immune from such lower bounds; slow decay can be established even for nonlinear benchmarks such as the stable width and entropy for high-dimensional parametric transport equations [15, 73]. Also relevant in this context are studies of stability of classical reduced models [32, 58, 21, 1].
Various attempts have been made to obtain the reduced complexity for convection-dominated problems by incorporating the structure of the transport. To provide a possibly incomplete list of proposed approaches that are the most relevant to this work: template-fitting [74], method of freezing [4, 59], shock reconstruction [16], approximated Lax-Pairs [28], advection modes [40], transported snapshot interpolation (TSI) [83], shifted proper orthogonal decomposition (sPOD) [67], calibrated manifolds [9], Lagrangian basis method [54], transport reversal [70], registration methods [78], Wasserstein barycenters [24] or quadratic manifolds [26]. Some of these proposed methods can be viewed as dynamical low-rank approximations [43, 75].
@@
The trained LRNRs can serve as a reduced model approximation of the family of solutions to the parametrized PDEs: A rapid approximation is achieved due to the small degrees of freedom, much like the rapid computations performed during the online phase for classical reduced models. Moreover, it is not difficult to see that a reduction in online complexity can be achieved by utilizing the low rank structure in the weights and biases,  [^1] During the review of this work, a concrete result demonstrating this observation has appeared in [12]. much like the classical reduced models [38]. This is a property LRNR inherits by being a direct generalization of the compositional reduced model utilizing transported subspaces [71].
@@
Organization of the Paper
The subject of this paper spans three different mathematical topics: The PDE theory of scalar conservation laws, reduced models, and neural networks. This paper is organized to cover all three aspects while aiming to keep each section as self-contained as possible.
  • Section 2. This section introduces new representation of the entropy solutions. The PDE theoretical presentation is self-contained and covers the entropy solution in its full generality; that is, solutions with arbitrary shock and rarefaction waves.
  • Section 3. A new neural network architecture called LRNR is introduced, inspired by the form of the new entropy solution representation from the previous section.
  • Section 4. We discuss the close relation between a subclass of LRNRs and the compositional reduced model called transported subspaces. We develop an important technique called the inverse-bias trick, and we prove that classical solutions to scalar conservation laws can be approximated efficiently by a LRNR with two layers and rank three. This section also serves as an introduction of new ideas with relatively light technical details.
  • Section 5. We develop a LRNR approximation to the entropy solutions by building on the new representation of the entropy solution developed in Section 2 and the inverse-bias trick from Section 4. We show that general entropy solutions can be approximated by a LRNR with five layers and rank two. This construction establishes a general low dimensional representation of the entropy solution, enabling an efficient representation of arbitrary shock wave interactions.
This paper introduces a substantial amount of new formulations and definitions. To aid the readers, the definitions and constructions are listed in a glossary in Appendix A. We will refer the reader to the glossary throughout the paper. Moreover, for brevity of presentation, the proofs of lemmas and theorems are postponed to Appendix B.
@@ 2.1 Entropy solutions
Characteristic curves of the initial value problem (2.1) are given by the solutions $X:\Omega_{x}\times\Omega_{t}\to\mathbb{R}$ to the family of ordinary differential equations
@@
This formulation for the characteristic curves is also called generalized characteristics [19]. It is straightforward to see that the problem (2.5) has a unique solution. This is due to the following: The RHS $G$ is locally continuous, so the unique solution exists locally where it is continuous [33]; the solution is uniformly stable in neighborhoods of the points of discontinuity due to the entropy conditions, and at points of discontinuity the shock curve is also uniquely determined by the Rankine-Hugoniot jump conditions [76]. One also observes that the unique solution $X$ is a continuous function of the initial value $x$, as $G$ is of bounded variation.
@@
We refer to both functions (2.5) and (2.6) as $G$ and distinguish them by the number of the independent variables (i.e. number of input arguments).
Since our problem has no source terms (2.1), it is constant along the characteristic curves. So we may rewrite (2.5) as follows,
@@ 2.2 Rarefied characteristics
In this section, we modify the differential equation describing the characteristic curves (2.7), so that the form of the entropy solution (2.11) can be simplified. The modified form of the solution combines the two cases into just one. We shall achieve this by extending the initial condition $u_{0}$ and the characteristic curves $X$ to an extended spatial domain that is larger than the original spatial domain $\Omega_{x}$.
@@
Figure 2: An example of an extended initial condition $\hat{u}_{0}$ of $u$. The newly assigned values of $\hat{u}_{0}$ in the interval $\Gamma_{1}$ are given by (2.20).
@@
and $\widehat{G}:\Omega_{x}\times\widehat{\Omega}_{x}\times\Omega_{t}\to\mathbb{R}$ is an extension of $G$ (2.6),
@@
The nonlinear RHS term (2.24) of the ODE (2.22) can be written as a function of bounded variation that depends only on $\widehat{X}$ and $t$. This implies that $\widehat{X}(\cdot,t)$ is non-decreasing for all $t\in\Omega_{t}$, regardless of whether $u_{0}$ has jumps (see, e.g. Ch. 5 of [33]); positive jumps of $u_{0}$ are replaced by continuous pieces in $\widehat{\Omega}_{x}$, negative jumps are replaced by piecewise constants.
See Figure 3 for an illustration showing $\hat{I}$ and its role in relating $u_{0}$ with $\hat{u}_{0}$ (2.22). A diagram depicting the evolution of $\widehat{X}$ is shown in Figure 4.
Figure 3: An example of the extension $u_{0}$ to $\hat{u}_{0}$ as in (2.20) for the Burgers flux $F(u)=u^{2}/2$. The function $\hat{I}^{+}$ has a jump discontinuity of magnitude $|\Gamma_{1}|$ as depicted in the upper right plot. Due to this jump, the values of $\hat{u}_{0}$ inside interval $\Gamma_{1}$ highlighted in the upper left plot are effectively omitted in the composed function $\hat{u}_{0}\circ\hat{I}^{+}$. This effect is shown in the lower right plot.
@@
The entropy solution $u$ to (2.1) is given by
@@
Recall that $(\cdot)^{+}$ denotes the left-inverse (2.10).
@@
See Appendix B.1.
We make a couple of remarks here. Note that new shocks cannot form at $\widehat{X}(x_{0},t)$ for $x_{0}\in\Gamma_{x}$ since the solution to (2.22) is strictly increasing with respect to $x$ at $x_{0}$. So the solution given by (2.25) at the points $(\widehat{X}(x_{0},t),t)$, for times before $\widehat{X}(x_{0},\cdot)$ enters a shock, agrees with the expression for the rarefaction fan. These observations imply that $\widehat{X}$ is continuous.
@@ 2.4 Compositional form
In the previous section, we have simplified the expression of the entropy solution from (2.11) to (B.1), by incorporating the evolution of the rarefaction waves into the characteristic curves in (2.22). Next, we simplify how the shock is represented in $\widehat{X}$.
In the classical regime, the characteristic curves (2.5) are represented by a linear superposition,
@@
Figure 5: A stationary shock example. Due to the jump discontinuity in the initial condition $\hat{u}_{0}$ located at $x=0.5$, the shock forms immediately at initial time. This shock is represented as the constant part of the function $\widehat{X}$ in $\Upsilon_{t}$, and the left-inverse of the rarefied characteristic $\widehat{X}^{+}$ has a jump discontinuity there. As time evolves, the characteristics merge into the shock, so $\Upsilon_{t}$ where $\widehat{X}$ is constant correspondingly expands over time. Consequently, the kinks in $\widehat{X}$ at the left and right endpoints of $\Upsilon_{t}$ travels to the left and to the right, respectively. (For a plot of the evolution of the rarefied characteristics, see Figure 7 below.)
@@
In this representation, the time-dependence is in the coefficients of linear combinations within each composition, and the dependence of the coefficients with respect to $t$ is linear, in a similar manner as the classical case. To elaborate, the first linear combination in (2.34) is $t\varphi_{1}-\varphi_{2}$ with two coefficient values $t$ and $-1$. Only the first coefficient depends on $t$ and the dependence is linear.
@@
We refer the reader to Appendix A for a glossary of various entropy solution formulations introduced in this section.
Figure 6: A diagram depicting the relief characteristics $\widecheck{X}$ after shock formation time $t_{S}$ is shown, along with the rarefied characteristics with multivalued inverse $\widehat{X}_{0}$. The relief characteristics use a different representation of the shock compared to the relief characteristics $\widehat{X}$, one that uses the addition of a box function $1_{\Upsilon_{t}}$. One observes that left-inverse $\widecheck{X}^{+}$ has a jump at $s(t)$, and this jump is identical to the corresponding one for $\widehat{X}^{+}$.
@@ 3 Low Rank Neural Representation
We introduce a family of feedforward neural networks that satisfy a low rank condition. The condition is inspired by the fact that the entropy solution, represented using the relief characteristic curves (2.34), is made up of compositions of low rank representations. To make this point clear, we draw up a diagram of each step of the evaluation of (2.34) along with the associated linear spaces, as follows.
@@
For example, the operations in the second step can be written as
@@
where $z=(x,t\varphi_{1}(x)-\varphi_{2}(x))^{T}$ is the output from the previous step. This expression is superposition of a few vectors, if one views the function values $\phi_{i}(z)$ as column vectors of infinite length. In addition, note that the individual the coefficients $[1,t,1]^{T}$ has either a linear dependence on $t$ or is constant.
This structure of repeated compositions closely resembles the repeated compositions in a neural network. The component that requires a close scrutiny is the left-inverse $(\cdot)^{+}$ and its treatment will be a key discussion in the coming sections (Sections 4.1 and 5.1). For the moment, let us focus on establishing an analogy with feedforward neural networks. We begin by introducing notations and definitions of a neural network. Our notations are similar to those in common use (see, e.g. [22, 3], for similar definitions). Throughout, the superscript $\ell$ will index the layer number.
@@
in which $\odot$ is the entrywise composition, $\sigma$ is the rectified linear unit (ReLU) $\sigma(x)=\max\{0,x\}$, and the affine maps $A_{\ell}$ are given by
@@
In what follows, we make the observation that the superpositions in between the compositions in (3.1) is analogous to the affine mappings that appear in the neural network (3.6). With this viewpoint, we show it is natural to consider a neural network with the affine mappings with low rank weights and biases, in correspondence to the low rank representation in (3.1). Let us consider the composition of two functions which are individually in the span of $\{\phi_{i}\}_{i=1}^{r}$ and $\{\varphi_{i}\}_{i=1}^{r}$, with the coefficients $\{\gamma_{i}^{3}\}_{i=1}^{r}$ and $\{\gamma_{i}^{2}\}_{i=1}^{r}$, resp.,
(3.7) $\left(\sum_{i=1}^{r}\gamma^{3}_{i}\phi_{i}\right)\circ\left(\sum_{i=1}^{r}\gamma^{2}_{i}\varphi_{i}\right)(x),$
which is a general form of the first composition in (3.1). For the simple scenario $r=1$ and both $\phi_{1},\varphi_{1}:\mathbb{R}\to\mathbb{R}$ are scalar valued 1d functions representable by 2-layer NNs, their composition has the structure shown in black and grey below,
(3.8) $(\gamma_{1}^{3}\phi_{1})\circ(\gamma_{1}^{2}{\color[rgb]{.5,.5,.5}\varphi_{1}})(x)$
$=\gamma_{1}^{3}\begin{bmatrix}\cdots\cdots\end{bmatrix}\sigma\odot\,\Bigg{(}\underbrace{\gamma_{1}^{2}\begin{bmatrix}\vdots\\ \vdots\end{bmatrix}{\color[rgb]{.5,.5,.5}\begin{bmatrix}\cdots\cdots\end{bmatrix}}}_{W^{2}}{\color[rgb]{.5,.5,.5}\sigma\odot\,\Bigg{(}\begin{bmatrix}\vdots\\ \vdots\end{bmatrix}}\begin{bmatrix}x\end{bmatrix}{\color[rgb]{.5,.5,.5}+\begin{bmatrix}\vdots\\ \vdots\end{bmatrix}\Bigg{)}+}\begin{bmatrix}\vdots\\ \vdots\end{bmatrix}{\color[rgb]{.5,.5,.5}[\cdot]}\Bigg{)}+[\cdot].$
In particular, the weight matrix of the second layer is rank one:
@@
Likewise, multiple scalar functions $\{\phi_{i}\}_{i=1}^{r}$ and $\{\varphi_{i}\}_{i=1}^{r}$ lead one to feedforward neural networks (3.5) whose weights and biases (3.6) are formed as linear combinations of rank-1 matrices in each layer. Assuming $r\ll M$, we have arrived at a low-rank form of the weight matrix.
This derivation illustrates that, in looking for a low-dimensional description of entropy solutions, it is natural to turn our attention to neural networks whose weights $W$ and biases in $B$ are represented as linear combinations of a few rank-1 linear transformations. For example, we require that
(3.10) $W^{\ell}=\sum_{i=1}^{r}\gamma_{i}^{\ell}U_{i}^{\ell},\quad B^{\ell}=\sum_{i=1}^{r}\theta_{i}^{\ell}V_{i}^{\ell},\qquad\quad\left\{\begin{aligned} \gamma_{i}^{\ell},&\theta_{i}^{\ell}\in\mathbb{R},\quad\\ U_{i}^{\ell}&\in\mathbb{W}^{\ell},V_{i}^{\ell}\in\mathbb{B}^{\ell}\text{ are rank-1,}\\ \ell&=1,...\,,L.\end{aligned}\right.$
The coefficients $\{\gamma^{\ell}_{i},\theta^{\ell}_{i}\}$ correspond to the coefficients $\{1,t\}$ used in the linear combinations at each compositional step in (3.1). One notes that the latter coefficients depend on the parameter $t$ (the time variable), and the role of this parameter is different from either the input $x$ or the fixed weights and biases $\{U_{i}^{\ell},V_{i}^{\ell}\}$ prescribing the linear spaces the weights and biases belong to. It is therefore natural to view the coefficients $\{\gamma_{i}^{\ell},\theta_{i}^{\ell}\}$ as distinct external parameters; a specific set of coefficients $\{\gamma_{i}^{\ell},\theta_{i}^{\ell}\}$ correspond to a specific neural network belonging to the family of neural networks whose weights and biases satisfy (3.10). The dependence on $t$ in (3.1) would correspond to the dependence of the coefficients $\{\gamma_{i}^{\ell},\theta_{i}^{\ell}\}$ on $t$. On the other hand, the fixed weights and biases correspond to the functions in the low rank representation that do not depend on $t$, e.g. the functions $\{\varphi_{i},\phi_{j},\psi_{k}\}$ in (3.2).
@@
In summary, a LRNR is a family of feedforward networks that have weights and bias at each individual layer that belong to a fixed linear subspace of dimension at most $r$. The aim of the following sections is to establish the connection between LRNR (Definition 3.3) and the entropy solution (3.1). Throughout this work, our focus is on scalar problems in a single spatial dimension, so the input and output dimensions are one, that is, $M_{0}=M_{L}=1$.
We refer the reader to Appendix A for a glossary of various notations and specifications regarding LRNRs introduced in this section.